# SKEL-CF: Coarse-to-Fine Biomechanical Skeleton and Surface Mesh Recovery

### v4

@@ 3.2 SKEL-CF
Methods 3DPW [35] Human3.6M [9] MOYO [33] MOYO-HARD
Two-stage Approaches
HMR2.0 + SKEL fit [39, 6] 81.0 54.4 53.6 34.1 130.5 93.7 - -
CameraHMR + SKEL fit [39, 27] 70.4 41.8 - - 75.5 49.9 88.7 61.4
End to End Approaches
HSMR [39] 81.5 54.8 50.4 32.9 104.5 79.6 120.0 97.7
SKEL-CF (Ours) 61.5 38.7 39.0 31.2 85.0 51.4 90.0 61.5
Table 1: Quantitative comparison of skeleton and surface mesh recovery using the SKEL [15] human model on 3DPW [35], Human3.6M [9], and MOYO [33]. The proposed SKEL-CF delivers consistent accuracy gains over previous end-to-end approach. denotes results reported by [39], denotes fitting the outputs of [27] into the SKEL [15] model.
@@
As illustrated in Figure 2, SKEL-CF employs an encoder–decoder architecture with a coarse-to-fine strategy to estimate SKEL parameters from a single RGB image. We denote the target parameter set as $\boldsymbol{\Theta}=\{\boldsymbol{\theta},\boldsymbol{\beta},\boldsymbol{\pi}\}$, where $\boldsymbol{\theta}$ represents the human pose, $\boldsymbol{\beta}$ the body shape, and $\boldsymbol{\pi}$ the camera extrinsics.
Given an input image, the encoder first encodes it into visual tokens and, together with camera intrinsic information, extracts contextual features through standard self-attention.
Using these features, the encoder produces initial estimates $\boldsymbol{\Theta}_{0}=\{\boldsymbol{\theta}_{0},\boldsymbol{\beta}_{0},\boldsymbol{\pi}_{0}\}$. These initial predictions are then progressively refined through $T$ decoder layers using cross-attention with the encoded visual features as context. The final output of SKEL-CF is $\boldsymbol{\Theta}_{T}=\{\boldsymbol{\theta}_{T},\boldsymbol{\beta}_{T},\boldsymbol{\pi}_{T}\}$.
@@
$\mathcal{L}_{\text{skel}}=\underbrace{\lambda_{kp}\mathcal{L}_{\text{kp}}}_{\text{keypoint-level}}+\underbrace{\lambda_{\beta}\mathcal{L}_{\beta}+\lambda_{\theta}\mathcal{L}_{\theta}}_{\text{parameter-level}}.$ (1)
where $\lambda_{kp}$, $\lambda_{\beta}$, and $\lambda_{\theta}$ are hyperparameters controlling the relative weight of each term. This formulation allows us to simultaneously supervise the reconstructed 3D/2D joints and the underlying pose and shape parameters. Note that no direct camera extrinsic supervision $\boldsymbol{\hat{\pi}}$ is provided; the ability to predict camera extrinsics is only implicitly learned through weak 2D keypoint reprojection, making the training more challenging.
@@
Since the proposed dataset 4DHuman-SKEL provides reliable pose $\hat{\boldsymbol{\theta}}$ and shape $\hat{\boldsymbol{\beta}}$ parameters, we impose supervision directly on these values:
@@
Figure 3: Visual comparison between the proposed SKEL-CF and HSMR [39]. Our proposed SKEL-CF achieves more precise skeletal estimations (best viewed in zoom). Additional visual results are provided in the supplementary material.
@@
SKEL-CF adopts a coarse-to-fine refinement strategy. The encoder first produces an initial estimate of the parameters, denoted as $\boldsymbol{\Theta}_{0}=\{\boldsymbol{\theta}_{0},\boldsymbol{\beta}_{0},\boldsymbol{\pi}_{0}\}$. The decoder then refines these predictions across $L$ layers, where the $i$-th layer outputs $\boldsymbol{\Theta}_{i}=\{\boldsymbol{\theta}_{i},\boldsymbol{\beta}_{i},\boldsymbol{\pi}_{i}\}$. The final layer produces the refined prediction $\boldsymbol{\Theta}_{L}$. Both the encoder output and the final decoder output are optimized using the same objective function $\mathcal{L}_{\text{skel}}$:
$\mathcal{L}_{\text{enc}}=\mathcal{L}_{\text{skel}},\quad\mathcal{L}_{\text{dec}}=\mathcal{L}_{\text{skel}}.$ (4)
@@
To stabilize training, we adopt an iterative refinement strategy with layer-wise supervision. Specifically, each decoder layer predicts a residual update to the parameters based on the output of the previous layer, progressively refining the estimate. To facilitate effective learning at each stage, we apply auxiliary supervision [44, 3] to every decoder layer using the same refinement loss.
$\mathcal{L}_{\text{refine}}=\sum_{i=1}^{L}\|\hat{\boldsymbol{\theta}}_{i}-\boldsymbol{\theta}_{i}\|_{1},$ (5)
where $L$ denotes the total number of decoder layers, and $\hat{\boldsymbol{\theta}}_{i}$ denotes the corresponding ground-truth pose parameters. We observe that optimizing $\mathcal{L}_{\text{skel}}$ provides similar performance. Since computing 2D keypoint predictions via the SKEL forward process is computationally expensive, we restrict optimization to the pose parameters.
Note that coarse-to-fine strategy, auxiliary supervision, and iterative refinement are widely adopted in the DETR-based object detection frameworks [44, 3]. The experiments to demonstrate the effectiveness of iterative refinement are provided in Appendix S2.
@@
The overall loss function combines the above components for both the coarse encoder prediction and the iterative refinement in decoder:
@@
Following HSMR [39], we adopt ViTPose-H [41], pretrained on COCO [22], as the encoder backbone. The hyper-parameters are set as $\lambda_{kp}=0.05$, $\lambda_{\beta}=0.0005$, $\lambda_{\theta}=0.001$, and $\lambda_{ref}=0.1$. More implementation details are provided in Appendix S3.
@@ 4 Experiment
Method 3DPW [35] EMDB [14] SPEC-SYN [17] MOYO [33] MOYO-HARD
[…]
SMPL [23]-based Approaches
SPEC [17] 96.5 53.2 118.5 138.9 87.7 161.3 83.5 56.9 98.9 - - - - - -
CLIFF [21] 69.0 43.0 81.2 103.5 68.3 123.7 128.5 55.8 139.0 154.6 109.3 155.7 - - -
HMR2.0a [6] 69.8 44.4 82.2 97.8 61.5 120.0 133.3 55.8 153.0 - - - - - -
TokenHMR [4] 70.5 43.8 86.0 88.1 49.8 104.2 110.5 51.8 127.6 - - - - - -
WHAM [30] 57.8 35.9 68.7 79.7 50.4 94.4 - - - - - - - - -
ReFit [36] 57.6 38.2 67.6 91.7 55.5 106.2 103.6 51.3 116.3 - - - - - -
CLIFF [21] 72.0 46.6 85.0 97.1 61.3 113.2 109.9 55.6 124.6 - - - - - -
HMR2.0b [6] 81.3 54.3 93.1 118.5 79.2 140.6 150.7 67.6 172.9 123.3 90.4 142.2 - - -
CameraHMR [27] 62.7 38.7 73.4 73.2 43.9 85.6 66.0 37.0 79.1 85.1 50.2 95.3 96.0 61.9 111.8
SKEL [15]-based Approaches
HSMR [39] 81.5 54.8 - - - - - - - 104.5 79.6 120.1 120.0 97.7 140.5
Table 2: Quantitative comparison of mesh recovery using the SMPL [23] human model on 3DPW [35], EMDB [14], SPEC-SYN [17], MOYO [33], and the more challenging MOYO-HARD dataset. The proposed SKEL-based method, SKEL-CF, achieves performance competitive with CameraHMR [27], with more pronounced improvements on complex motion scenarios as MOYO-HARD.
Figure 4: Visual comparison between the proposed SKEL-CF and CameraHMR [27]. The proposed SKEL-CF is built upon the SKEL [15] representation, which enforces anatomically consistent joint motion, resulting in more natural pose predictions compared to the SMPL [23]-based CameraHMR [27] (best viewed in zoom). Additional visual examples are provided in the supplementary material.
@@ 4.1 Datasets and evaluation metrics
We evaluate SKEL-CF on five representative datasets covering diverse environments and motion complexities, following standard practice [27, 39, 11, 4]. 3DPW [35] provides in-the-wild videos with accurate 3D annotations captured by moving cameras. Human3.6M [9] includes controlled indoor actions (e.g., sitting, walking, greeting) with motion-capture ground truth. EMDB [14] offers high-quality pose and shape sequences captured with electromagnetic sensors and handheld cameras. SPEC-SYN [17] is synthetic with explicit camera intrinsics and diverse perspectives for controlled camera evaluation. In contrast, MOYO [33] contains large-scale yoga sequences with extreme poses, frequent self-occlusions, and ground contact, providing the most challenging and diverse test scenario. Since most yoga videos in MOYO starts and ends with the same static poses, we select the front-view camera among the eight available viewpoints and remove the first and last 25% of frames to retain the complex and diverse motion segments, forming the curated MOYO-HARD subset. More details are provided in Appendix, Figure S1., where we additionally include visualizations of the removed first and last 25% segments of the MOYO videos.
@@
We compare SKEL-based methods in Table 1. Among end-to-end approaches, SKEL-CF consistently outperforms prior work HSMR [39] across all datasets. On the challenging MOYO [33] benchmark, SKEL-CF achieves up to 18.6% MPJPE and 35.4% PA-MPJPE improvements, demonstrating strong robustness under large pose variations and occlusions. We also evaluate two-stage baseline, HMR2.0 + SKEL fit and CameraHMR + SKEL fit. While CameraHMR + SKEL fit achieves competitive results on MOYO and MOYO-HARD, its performance drops on 3DPW, indicating limited stability across datasets. Moreover, the fitting-based approach is computationally expensive (approximately 4 min/img on RTX 3090), whereas SKEL-CF runs in a single forward pass (approximately 0.8 s/img), making it both more efficient and more reliable.
@@
We provide a visual comparison with HSMR [39] in Figure 3. As shown, the proposed SKEL-CF achieves more accurate skeletal estimation and mesh reconstruction. More visual examples similar to Figure 3 are included in Appendix, Figure S3.
Figure 5: Illustration of the mesh refinement. The light green meshes represent the initial coarse estimations, while the dark green meshes denote the progressively refined final results.
@@ 4.3 Comparison with SMPL-based approaches
We further compare SKEL-CF with recent state-of-the-art SMPL-based human mesh recovery methods across four representative datasets: 3DPW [35], EMDB [14], SPEC-SYN [17], MOYO [33] and more challenging MOYO-HARD, as summarized in Table 2. Note that SMPL [23] is less constrained compared to SKEL [15], enable achieving strong quantitative performance but occasionally producing anatomical inconsistency or unnatural joint motions(see Figure 1). Despite this inherent difference, SKEL-CF achieves comparable performance to CameraHMR [27], the strongest SMPL-based model trained on 4DHuman, demonstrating that kinematically constrained SKEL representations can reach a similar level of numerical precision while maintaining stronger physical plausibility.
Notably, on the challenging MOYO [33] dataset, which contains complex yoga sequences with strong self-occlusion and body–ground interactions, SKEL-CF achieves state-of-the-art performance in dense mesh prediction (PVE) and overall reconstruction accuracy. Furthermore, on the MOYO-HARD subset, focused on dynamic and highly articulated poses, SKEL-CF consistently outperforms all SMPL-based baselines across all metrics (MPJPE, PA-MPJPE, and PVE). This highlights its superior ability to maintain accurate and physically realistic motion representations under extreme articulations. These results demonstrate that SKEL-CF provides a scalable and anatomically faithful framework for human motion analysis. For completeness, we also report the 2D keypoint results in Table 3. Our approach achieves comparable accuracy under PCK@0.1, indicating that the projected mesh provides a consistent and accurate fit to the 2D image plane.
@@
We provide a visual comparison with the SMPL [23]-based state-of-the-art method CameraHMR [27] in Figure 4. As shown, the proposed SKEL-CF produces anatomically more consistent joint motions, benefiting from the structural constraints of the SKEL [15] model. In contrast, CameraHMR tends to generate unnatural joint configurations under complex motion scenarios. Additional visualizations similar to Figure 4 are provided in Appendix, Figure S2.

### v5

@@ 3.2 SKEL-CF
As illustrated in Figure 2, SKEL-CF adopts an encoder–decoder architecture with a coarse-to-fine strategy to estimate SKEL parameters from a single RGB image. The target parameter set is defined as $\boldsymbol{\Theta}=\{\boldsymbol{\theta},\boldsymbol{\beta},\boldsymbol{\pi}\}$, where $\boldsymbol{\theta}$ denotes human pose, $\boldsymbol{\beta}$ is body shape, and $\boldsymbol{\pi}$ indicates camera extrinsics. Given an input image, following previous work [patel2024camerahmr], we first detect the person using an off-the-shelf human detector and convert the cropped region into visual tokens. We also estimate the camera intrinsic parameter, specifically the focal length $f$, from the full image using a pretrained predictor from CameraHMR [patel2024camerahmr], which remains frozen during training. The visual tokens are passed through the encoder to produce initial predictions $\boldsymbol{\Theta}_{0}={\boldsymbol{\theta}_{0},\boldsymbol{\beta}_{0},\boldsymbol{\pi}0}$ along with contextual visual features. These predictions are then progressively refined through $T$ decoder layers. At each stage, refinement is guided by the encoded visual features and geometric cues $\mathcal{F}_{\text{bbox}}=\left(\frac{c_{x}}{f},\frac{c_{y}}{f},\frac{s}{f}\right)$, where $(c_{x},c_{y})$ and $s$ denote the bounding box center and scale. After $T$ refinement steps, the final output is $\boldsymbol{\Theta}_{T}=\{\boldsymbol{\theta}_{T},\boldsymbol{\beta}_{T},\boldsymbol{\pi}_{T}\}$, yielding progressively improved SKEL parameter estimates.
@@
$\mathcal{L}_{\text{skel}}(\boldsymbol{\Theta},\boldsymbol{\hat{\Theta}})=\underbrace{\lambda_{kp}\mathcal{L}_{\text{kp}}}_{\text{keypoint-level}}+\underbrace{\lambda_{\beta}\mathcal{L}_{\beta}+\lambda_{\theta}\mathcal{L}_{\theta}}_{\text{parameter-level}}.$ (1)
where $\lambda_{kp}$, $\lambda_{\beta}$, and $\lambda_{\theta}$ are hyperparameters controlling the relative weight of each term. This formulation allows us to simultaneously supervise the reconstructed 3D/2D joints and the underlying pose and shape parameters. Note that no direct camera extrinsic supervision $\boldsymbol{\hat{\pi}}$ is provided; the ability to predict camera extrinsics is only implicitly learned through 2D keypoint reprojection, making the training challenging.
@@
Since the proposed dataset HMR-SKEL provides reliable pose $\hat{\boldsymbol{\theta}}$ and shape $\hat{\boldsymbol{\beta}}$ parameters, we impose supervision directly on these values:
@@
SKEL-CF adopts a coarse-to-fine refinement strategy. The encoder first produces an initial estimate of the parameters, denoted as $\boldsymbol{\Theta}_{0}=\{\boldsymbol{\theta}_{0},\boldsymbol{\beta}_{0},\boldsymbol{\pi}_{0}\}$. The decoder then refines these predictions across $T$ layers, where the $i$-th layer outputs $\boldsymbol{\Theta}_{i}=\{\boldsymbol{\theta}_{i},\boldsymbol{\beta}_{i},\boldsymbol{\pi}_{i}\}$. The final layer produces the refined prediction $\boldsymbol{\Theta}_{T}$. Both the encoder output and the final decoder output are optimized using the same objective function $\mathcal{L}_{\text{skel}}$:
[…]
@@
To stabilize training, we adopt an iterative refinement strategy with layer-wise supervision. Specifically, each decoder layer predicts a residual update to the parameters based on the output of the previous layer, iteratively refining the estimate. To facilitate effective learning at each stage, we apply auxiliary supervision [zhu2020deformable, detr] to intermediate decoder layers using the same refinement loss.
$\mathcal{L}_{\text{refine}}=\sum_{i=1}^{T-1}\|\hat{\boldsymbol{\theta}}_{i}-\boldsymbol{\theta}_{i}\|_{1},$ (5)
where $T$ denotes the total number of decoder layers, and $\hat{\boldsymbol{\theta}}_{i}$ denotes the corresponding ground-truth pose parameters. We observe that optimizing $\mathcal{L}_{\text{skel}}$ at intermediate decoder layers achieves performance comparable to $\mathcal{L}_{\text{refine}}$. However, as computing keypoint-level predictions through the SKEL forward process is computationally expensive, we limit the optimization to the pose parameters for efficiency.
Note that coarse-to-fine strategy and iterative refinement are widely adopted in the DETR-based object detection frameworks [zhu2020deformable, detr]. The experiments to demonstrate the effectiveness of iterative refinement are provided in the supplementary material.
@@
The total training objective $\mathcal{L}_{\text{tot}}$ aggregates the encoder and decoder losses, together with the refinement term:
@@
Following HSMR [xia2025hsmr], we adopt ViTPose-H [xu2022vitpose], pretrained on COCO [coco], as the encoder backbone. The hyper-parameters are set as $\lambda_{kp}=0.05$, $\lambda_{\beta}=0.0005$, $\lambda_{\theta}=0.001$, and $\lambda_{ref}=0.1$. More implementation details are provided in the supplementary material.
@@ 4.1 Datasets and evaluation metrics
We evaluate SKEL-CF on five representative datasets covering diverse environments and motion complexities, following standard practice [patel2024camerahmr, xia2025hsmr, kanazawa2018hmr, dwivedi2024tokenhmr]. 3DPW [3dpw] provides in-the-wild videos with accurate 3D annotations captured by moving cameras. Human3.6M [human36m] includes controlled indoor actions (e.g., sitting, walking, greeting) with motion-capture ground truth. EMDB [kaufmann2023emdb] offers high-quality pose and shape sequences captured with electromagnetic sensors and handheld cameras. SPEC-SYN [spec-syn] is a synthetic dataset with diverse camera variations, enabling controlled evaluation under varying camera settings. In contrast, MOYO [moyo] contains large-scale yoga sequences with extreme poses, frequent self-occlusions, and ground contact, providing the most challenging and diverse test scenario.
Since most yoga videos in MOYO starts and ends with the same static poses, we select the front-view camera among the eight available viewpoints and remove the first and last 25% of frames to retain the complex and diverse motion segments, forming the curated MOYO-HARD subset. More details about MOYO-HARD are provided in the supplementary material, where we include visualizations of the removed first and last 25% segments of the MOYO videos.
@@
Table 1: Quantitative comparison of skeleton and surface mesh recovery using the SKEL [SKEL] human model on 3DPW [3dpw], Human3.6M [human36m], and MOYO [moyo]. The proposed SKEL-CF delivers consistent accuracy gains over previous end-to-end approach. denotes results reported by [xia2025hsmr], denotes fitting the outputs of [patel2024camerahmr] to the SKEL [SKEL] model.
Methods 3DPW [3dpw] Human3.6M [human36m] MOYO [moyo] MOYO-HARD
Two-stage Approaches
HMR2.0 + SKEL fit [xia2025hsmr, goel20234dhuman] 81.0 54.4 53.6 34.1 130.5 93.7 - -
CameraHMR + SKEL fit [xia2025hsmr, patel2024camerahmr] 70.4 41.8 - - 75.5 49.9 88.7 61.4
End-to-end Approaches
HSMR [xia2025hsmr] 81.5 54.8 50.4 32.9 104.5 79.6 120.0 97.7
SKEL-CF (Ours) 61.5 38.7 39.0 31.2 85.0 51.4 90.0 61.5
Figure 3: Visual comparison between the proposed SKEL-CF and HSMR [xia2025hsmr]. Our proposed SKEL-CF achieves more precise skeletal estimations (best viewed in zoom). Additional visual results are provided in the supplementary material.
@@
We compare SKEL-based methods in Table 1. Among end-to-end approaches, SKEL-CF consistently outperforms prior work HSMR [xia2025hsmr] across all datasets. On the challenging MOYO [moyo] benchmark, SKEL-CF achieves up to 18.6% MPJPE and 35.4% PA-MPJPE improvements, demonstrating strong robustness under large pose variations and occlusions. We also evaluate two two-stage baselines that first estimate SMPL parameters and then fit them to SKEL: HMR2.0 + SKEL fitting and CameraHMR + SKEL fitting. While CameraHMR + SKEL fit achieves competitive results on MOYO and MOYO-HARD, its performance drops on 3DPW, indicating limited stability across datasets. Moreover, the fitting-based approach is computationally expensive (approximately 4 minute per image on RTX 3090), whereas SKEL-CF runs in a single forward pass (approximately 0.8 second per image), making it both more efficient and more reliable.
@@
We provide a visual comparison with HSMR [xia2025hsmr] in Figure 3. As shown, the proposed SKEL-CF achieves more accurate skeleton and surface reconstruction. More visual examples similar to Figure 3 are included in supplementary material.
Table 2: Quantitative comparison of mesh recovery using the SMPL [loper2015smpl] human model on 3DPW [3dpw], EMDB [kaufmann2023emdb], SPEC-SYN [spec-syn], MOYO [moyo], and the more challenging MOYO-HARD dataset. The proposed SKEL-based method, SKEL-CF, achieves performance competitive with CameraHMR [patel2024camerahmr], with more pronounced improvements on complex motion scenarios as MOYO-HARD.
Method 3DPW [3dpw] EMDB [kaufmann2023emdb] SPEC-SYN [spec-syn] MOYO [moyo] MOYO-HARD
[…]
SMPL [loper2015smpl]-based Approaches
SPEC [spec-syn] 96.5 53.2 118.5 138.9 87.7 161.3 83.5 56.9 98.9 - - - - - -
CLIFF [li2022cliff] 69.0 43.0 81.2 103.5 68.3 123.7 128.5 55.8 139.0 154.6 109.3 155.7 - - -
HMR2.0a [goel20234dhuman] 69.8 44.4 82.2 97.8 61.5 120.0 133.3 55.8 153.0 - - - - - -
TokenHMR [dwivedi2024tokenhmr] 70.5 43.8 86.0 88.1 49.8 104.2 110.5 51.8 127.6 - - - - - -
WHAM [shin2024wham] 57.8 35.9 68.7 79.7 50.4 94.4 - - - - - - - - -
ReFit [wang2023refit] 57.6 38.2 67.6 91.7 55.5 106.2 103.6 51.3 116.3 - - - - - -
CLIFF [li2022cliff] 72.0 46.6 85.0 97.1 61.3 113.2 109.9 55.6 124.6 - - - - - -
HMR2.0b [goel20234dhuman] 81.3 54.3 93.1 118.5 79.2 140.6 150.7 67.6 172.9 123.3 90.4 142.2 - - -
CameraHMR [patel2024camerahmr] 62.7 38.7 73.4 73.2 43.9 85.6 66.0 37.0 79.1 85.1 50.2 95.3 96.0 61.9 111.8
SKEL [SKEL]-based Approaches
HSMR [xia2025hsmr] 81.5 54.8 - - - - - - - 104.5 79.6 120.1 120.0 97.7 140.5
Figure 4: Visual comparison between the proposed SKEL-CF and CameraHMR [patel2024camerahmr]. The proposed SKEL-CF is built upon the SKEL [SKEL] representation, which enforces anatomically consistent joint motion, resulting in more natural pose predictions compared to the SMPL [loper2015smpl]-based CameraHMR [patel2024camerahmr] (best viewed in zoom). Additional visual examples are provided in the supplementary material.
@@ 4.3 Comparison with SMPL-based approaches
We further compare SKEL-CF with recent state-of-the-art SMPL-based human mesh recovery methods across four representative datasets: 3DPW [3dpw], EMDB [kaufmann2023emdb], SPEC-SYN [spec-syn], MOYO [moyo] and more challenging MOYO-HARD, as summarized in Table 2. Note that SMPL [loper2015smpl] is less constrained compared to SKEL [SKEL], enabling strong quantitative performance but occasionally producing anatomical inconsistency or unnatural joint motions(see Figure 1). Despite this inherent difference, SKEL-CF achieves comparable performance to CameraHMR [patel2024camerahmr], the strongest SMPL-based model trained on 4DHuman dataset, demonstrating that kinematically constrained SKEL representations can reach a similar level of numerical precision while maintaining stronger physical plausibility.
Notably, on the challenging MOYO [moyo] dataset, which contains complex yoga sequences with strong self-occlusion and body–ground interactions, SKEL-CF achieves state-of-the-art performance in dense mesh prediction (PVE) and overall reconstruction accuracy. Furthermore, on the MOYO-HARD subset, focused on dynamic and highly articulated poses, SKEL-CF consistently outperforms all SMPL-based baselines across all metrics (MPJPE, PA-MPJPE, and PVE). This highlights its superior ability to maintain accurate and physically realistic motion representations under extreme articulations. These results demonstrate that SKEL-CF provides a scalable and anatomically faithful framework for human motion analysis. For completeness, we also report the 2D keypoint results in Table 3. Our approach achieves comparable accuracy under PCK@0.1, indicating that the projected mesh provides a consistent and accurate fit to the 2D image plane.
Table 3: 2D Keypoint Performance on COCO [coco]. For completeness, we also report PCK scores of the projected keypoints at thresholds of 0.05 and 0.1, showing that SKEL-CF achieves comparable 2D keypoint quality while maintaining significantly better biomechanical mesh quality.
Method PCK@0.05 $\uparrow$ PCK@0.1 $\uparrow$
PyMAF [pymaf2021] 0.68 0.86
CLIFF [li2022cliff] 0.64 0.88
PARE [kocabas2021pare] 0.72 0.91
PyMAF-X [pymafx2023] 0.79 0.93
HMR2.0a [goel20234dhuman] 0.79 0.94
CameraHMR [patel2024camerahmr] 0.84 0.94
HSMR [xia2025hsmr] 0.85 0.96
SKEL-CF (Ours) 0.80 0.93
@@
We provide a visual comparison with the SMPL [loper2015smpl]-based state-of-the-art method CameraHMR [patel2024camerahmr] in Figure 4. As shown, the proposed SKEL-CF produces anatomically more consistent joint motions, benefiting from the structural constraints of the SKEL [SKEL] model. In contrast, CameraHMR tends to generate unnatural joint configurations under complex motion scenarios. Additional visualizations similar to Figure 4 are provided in supplementary material.